# Ranked Set Sampling-Based Multilayer Perceptron: Improving Generalization via Variance-Based Bounds

### v1

@@ 1 Introduction
The evaluation of the generalization ability of machine learning models is one of the core issues in statistical learning theory. The generalization ability of existing learning models is often analyzed through frameworks such as VC dimension  [33], Rademacher complexity [3], margin distribution [18, 20], stability analysis [5], and PAC-Bayes theory [32]. While these approaches are theoretically significant, they predominantly focus on the complexity of the overall hypothesis space while neglecting the crucial influence of data distribution on generalization. Traditional generalization bounds tend to be overly loose, failing to accurately reflect true model performance. Model variance, a critical metric for assessing prediction stability under specific data distributions, is directly linked to generalization capabilities. We extend existing theory by deriving a widely applicable error bound that establishes the connection between empirical loss variance and generalization. Our findings demonstrate that decreasing variance can substantially enhance model performance.
@@ 3 A Variance-Based Generalization Bound
Usually, the underlying probability distribution of $\mathcal{X}\times\mathcal{Y}$ is unknown. We merely have a set of independently empirical data drawn from it, denoted as $D=\{(\textbf{{x}}_{1},y_{1}),...,(\textbf{{x}}_{N},y_{N})\}$, where $N$ is the number of object. Through the training data set, classifiers can be obtained by optimizing the empirical convex risk functions: $\hat{\mathit{R}}_{\phi}(f)=\frac{1}{N}\sum_{i=1}^{N}\phi(y_{i},f(\textit{{x}}_{i}))$
@@ 13 CIFAR-10 60000 32×32×3 10
Evaluation Method and Metrics. For each dataset, we conduct a random partition into training and testing sets with a ratio of $7$:$3$. This partition is executed 30 times to obtain a comparison of average performance. We employ accuracy and $F_{1}$-score to evaluate the performance of each method. The larger values indicate a better classification performance.
@@
Implementation Details. The weak classifier used is a network with three fully connected layers, whose structure is shown in Figure 1. Each layer of the network consists of a linear layer, a batch normalization layer, and a ReLU activation layer. The output dimensions of the three layers are set to 256, 128 and the number of classes. The number of base classifiers $T$ is set to $T=51$. The $K$ in Eq. (1) is set as $\lfloor\sqrt{N}\rfloor$, which is the maximum possible number of classes for N samples [28]. The batch size number is set to $32$.
@@ 7.2 Proof of Theorem 2
Theorem 2. (Equal Expectation) For the additive convex loss function $\phi$ and an independent identically distribution sample, the expectations of empirical risks based on the SRS sample and RSS sample are equal: $\mathbb{E}\hat{R}_{\phi,RSS}(f)=\mathbb{E}\hat{R}_{\phi,SRS}(f)$
@@
[…]
@@ 7.3 Proof of Theorem 3
where EulerE is the Euler function. When $n$ is an odd number larger than $3$, $EulerE[n-1,0]=0$. Thus, we get $b_{\phi}=-\frac{1}{2}$. Then, Eq. (19) can be rewritten as follows:

### v2

@@ 1 Introduction
The evaluation of the generalization ability of machine learning models is one of the core issues in statistical learning theory. The generalization ability of existing learning models is often analyzed through frameworks such as VC dimension  [37], Rademacher complexity [3, 50], margin distribution [18, 20], stability analysis [5, 28], and PAC-Bayes theory [36]. While these approaches are theoretically significant, they predominantly focus on the complexity of the overall hypothesis space while neglecting the crucial influence of data distribution on generalization. Traditional generalization bounds tend to be overly loose, failing to accurately reflect true model performance. Model variance, a critical metric for assessing prediction stability under specific data distributions, is directly linked to generalization capabilities. We extend existing theory by deriving a widely applicable error bound that establishes the connection between empirical loss variance and generalization. Our findings demonstrate that decreasing variance can substantially enhance model performance.
@@ 3 A Variance-Based Generalization Bound
Usually, the underlying probability distribution of $\mathcal{X}\times\mathcal{Y}$ is unknown. We merely have a set of independently empirical data drawn from it, denoted as $D=\{(\textbf{{x}}_{1},y_{1}),...,(\textbf{{x}}_{N},y_{N})\}$, where $N$ is the number of object. Through the training data set, classifiers can be obtained by optimizing the empirical convex risk functions: $\hat{\mathit{R}}_{\phi}(f)=\frac{1}{N}\sum_{i=1}^{N}\phi(y_{i},f(\textit{{x}}_{i}))$ .
@@ 13 CIFAR-10 60000 32×32×3 10
Evaluation Method and Metrics. For each dataset, we conduct a random partition into training and testing sets with a ratio of 77:33. This partition is executed 30 times to obtain a comparison of average performance. We employ accuracy and $F_{1}$-score to evaluate the performance of each method. The larger values indicate a better classification performance.
@@
Implementation Details. The weak classifier used is a network with three fully connected layers, whose structure is shown in Figure 1. Each layer of the network consists of a linear layer, a batch normalization layer, and a ReLU activation layer. The output dimensions of the three layers are set to 256, 128 and the number of classes. The number of base classifiers $T$ is set to $T=51$. The $K$ in Eq. (1) is set as $\lfloor\sqrt{N}\rfloor$, which is the maximum possible number of classes for N samples [32, 31, 30, 29]. The batch size number is set to $32$.
@@ 7.2 Proof of Theorem 2
Theorem 2. (Equal Expectation) For the additive convex loss function $\phi$ and an independent identically distribution sample, the expectations of empirical risks based on the SRS sample and RSS sample are equal: $\mathbb{E}\hat{R}_{\phi,RSS}(f)=\mathbb{E}\hat{R}_{\phi,SRS}(f)$ .
@@
$=\frac{1}{K}\sum_{r=1}^{K}\int\phi(y,f(\textit{{x}}))g_{[r]}(f(\textit{{x}}),y)d\textit{{x}}=\int\phi(y,f(\textit{{x}}))g(f(\textit{{x}}),y)d\textit{{x}}$
@@ 7.3 Proof of Theorem 3
where EulerE is the Euler function. When $n$ is an odd number larger than $3$, EulerE[n1,0]=0EulerE[n-1,0]=0. Thus, we get $b_{\phi}=-\frac{1}{2}$. Then, Eq. (19) can be rewritten as follows: